### 20 Mar 2026 v4

# Algebraic Characterizations of Angle Multisections over Rings

Abstract
Let be integers, and let be a subring of with field of fractions In this article, we generalize the rational angle bisection problem previously proposed by the author as follows: characterize the pairs of linearly independent vectors that form angles with sequences of -sector vectors lying in When and are nonorthogonal, we prove that this condition is equivalent to the existence of a root in of a certain -th degree polynomial over In particular, when the condition holds if and only if the polynomial has a root among the divisors of its constant term. When with integer we also prove that the condition is equivalent to where is the angle between and
1 Introduction
Throughout this article, let be integers, and let be a subring of with field of fractions For any two nonzero vectors we denote by the angle between and as well as its measure in depending on the context.
Definition 1.
Let
A nonzero vector is called an -th -sector vector of over (or simply, an angle multisector vector over ), if there exist a sequence of nonzero coplanar vectors in and an integer such that for some and
for each In this case, we also say that the vectors form a sequence of -equisector vectors of over (or simply, angle equisector vectors over ).
The angle between and is said to be -sectable over if there exists a sequence of -equisector vectors of over
In this article, we generalize the rational angle bisection problem previously proposed by the author (see [2, Problem 1]) to the following problem, which may be referred to as the rational angle multisection problem.
Problem 1.
Characterize the pairs of linearly independent vectors that form -sectable angles over
Remark 1.
Let with and Then is -sectable over if and only if is -sectable over
We utilize the following proposition derived from the Householder reflection (see [4, Section 1.1]). For any two vectors we denote their inner product by
Proposition 1.
Let be linearly independent vectors. Then there exists a nonzero vector such that is a bisector vector of Such a vector is parallel to
Proposition 1 is fundamental and plays an important role in practical applications, as it enables us to extend a sequence of angle equisector vectors over to an arbitrary length (see Section 2).
Example 1.
Let and Then
and therefore has a sequence of nonasector vectors. That is, in the -plane, the obtuse angle between the two lines
has a sequence of nonasector lines
as in Figure 1.
The following theorem is known for the case when Let
Theorem 1 ([2, Theorem 1]).
Let be linearly independent vectors. The following conditions are equivalent.
A bisector vector of exists in
A normal vector of a bisecting hyperplane between the two hyperplanes with normal vectors and exists in
There exists a vector such that
We have
Remark 2.
If (A4) is satisfied, then in (A1)β(A3) is parallel to
where is a positive number in such that
In Section 3, we prove the following theorem.
Theorem 2.
Let be linearly independent and nonorthogonal vectors. Let
and
The following conditions are equivalent.
is -sectable over
We have
The -th degree polynomial
(1.3) in over has a root in
Furthermore, if (1.3) has a root then a first -sector vector of over is given by
for some
Combining Theorem 2 with the rational root theorem, we obtain the following corollary.
Corollary 1.
Let be an algebraic number field with class number and let be its ring of integers. Let be linearly independent and nonorthogonal vectors. Then is -sectable over if and only if (1.3) has a root in
By the rational root theorem, if then the possible integral roots of (1.3) are restricted to the divisors of if is even, or those of if is odd.
Example 2.
Let Then is trisectable over since the cubic polynomial
corresponding to (1.5) has an integral root A first trisector vector of over is given by
Condition (A4) in Theorem 1 is equivalent to We generalize this fact to the following theorem in Section 4.
Theorem 3.
Let and let be an integer. The following conditions are equivalent.
is -sectable over
We have
Theorem 3 enables us to verify the -sectability of over through purely algebraic calculations. Specifically, (N2) can be verified by checking (A4) and
for each These conditions ensure that the geometric steps of the successive bisections over are equivalent to checking for the existence of certain square roots within
Example 3.
Let and Then is quadrisectable over since
and
2 Extending Sequences of Angle Equisectors
In this section, we provide a brief proof of Proposition 1 and state its corollaries.
Proof of Proposition 1.
It suffices to show that for
If then since
Suppose that Then
which implies Therefore, a bisector vector of is given by
which is parallel to Thus, we conclude that β
Proposition 1 enables us to construct a sequence of angle equisector vectors over by specifying the first two terms.
Example 4.
Let and Then
Let Then is a bisector vector of and
Let Then is a bisector vector of and
Let Then is a bisector vector of These imply that has a sequence of quadrisector vectors. That is, in the -space, the acute angle between the two lines
has a sequence of quadrisector lines
Proposition 1 immediately yields the following three corollaries.
Corollary 2.
A given sequence of angle equisector vectors over can be extended to an arbitrary length. Specifically, for any integers and such that any sequence of angle -equisector vectors over can be extended to a sequence of angle -equisector vectors over by appending suitable vectors in
Corollary 3.
For any nonzero vectors there exists an infinite sequence of nonzero vectors in such that
Corollary 4.
Let If has a first -sector vector in then is -sectable over
Corollary 5.
Let Let be pairwise coprime integers greater than and let Then the following conditions are equivalent.
is -sectable over
is -sectable over for each
Proof.
It is obvious that (i) implies (ii). To show that (ii) implies (i), it suffices to consider the case when Suppose that is both -sectable and -sectable over Since and are coprime, there exist integers such that Then a -th -sector vector and a -th -sector vector of can be taken in These two vectors form an angle of Thus, for each a -th -sector vector can be taken in by Corollary 3. This implies that (i) holds. β
3 Angle Multisection
In this section, we provide a proof and an example of Theorem 2. Our arguments are based on the following three lemmas.
Lemma 1.
Proof.
These relations follow directly from the definition of the inner product,
and the formula for the area of the parallelogram spanned by and
Lemma 2.
Proof.
Let Suppose that satisfies Since and are linearly independent and nonorthogonal, which implies and is well-defined. Let where Substituting
and
into we obtain
Let Substituting into (3.7), we obtain
Multiplying both sides by we find that is a root of (1.3). Since there exist distinct values of corresponding to the solutions of within we conclude that (1.3) has distinct nonzero real roots. β
Lemma 3.
Let be linearly independent vectors, and let be defined as in (1.2). If and and are nonorthogonal, then
Proof.
Let and Then there exist scalars not both zero, such that We use the orthonormal basis of defined in (3.1) and (3.2). On one hand, by (3.3), can be expressed as
On the other hand, substituting (3.4) into the relation , we have
Since and are linearly independent over we can equate the coefficients to obtain
Dividing the second equation by the first yields
which implies (3.8) since β
Proof of Theorem 2.
By Lemma 3, (M1) implies (M2).
Furthermore, (M2) implies (M3) as a direct consequence of Lemma 2.
Example 5.
Example 6.
Let Then is trisectable over since the cubic polynomial
corresponding to (1.5) has an integral root A first trisector vector of over is given by
4 Angle -section
In this section, we provide a proof and an example of Theorem 3.
Proof of Theorem 3.
Let
For the case when note that (A1) in Theorem 1 is equivalent to Indeed, by Theorem 1, (A1) is equivalent to (A4), which can be restated as or equivalently, since and
We now consider the general case when Suppose that (N1) is satisfied. In a sequence of -equisector vectors of the vector is a bisector vector of which implies (N2) by the above argument for Conversely, suppose that (N2) is satisfied. Then holds for each by the double-angle formula. First, implies that has a bisector vector Second, implies that has a bisector vector which serves as a first quadrisector vector of By repeating this process times, we find that has a first -sector vector Thus, is -sectable over by Corollary 4. β
Example 7.
Theorem 3 remains valid for the case when For example, if then and has a bisector vector which is consistent with the fact that implies (A4).
Example 8.
Let and Then is quadrisectable over since
and